OpenStudy (anonymous):

find the solution set 3 lessthanequalto 2x-1 lessthanequalto 11

OpenStudy (amistre64):

3 <= 2x -1 <= 11 ??

OpenStudy (amistre64):

add +1 to everything 3+1 <= 2x -1+1 <= 11+1 4 <= 2x <= 12 now divide the everything by 2 4/2 <= 2x/2 <= 12/2 2 <= x <= 6 x = [2,6]